# Hénon maps with many rational periodic points

### v1

@@ Abstract.
Building on work of Doyle and Hyde on polynomial maps in one variable, we produce for each odd integer $d\geq 2$ a Hénon map of degree $d$ defined over $\mathbb{Q}$ with at least $(d-4)^{2}$ integral periodic points. This provides a quadratic lower bound on any conjectural uniform bound for periodic rational points of Hénon maps. In contrast with the work of Doyle and Hyde, our examples also admit integer cycles of large period.
@@ 1. Introduction
In this article, we study the polynomials constructed by Doyle and Hyde and use them to construct an explicit family of Hénon maps $\{h_{d}(x,y)\}$ over $\mathbb{Q}$ with degree $d\rightarrow\infty$ and a large number of rational periodic points. This answers in the affirmative a question of Ingram [henon_openqs].
@@
A direct computation shows (see section 5.2) that any point of $\mathbb{C}^{2}$ with integer coordinates is periodic under iteration of the limiting Hénon map $h_{\infty}(x,y)=(y,-x+s_{\infty}(y)$, with period $n\in\{1,4,5,6,12,20\}$ determined by the class of $(x,y)$ modulo $6$. Though the proof of Theorem A is not asymptotic, each map $h_{d}$ preserves the lattice $\mathbb{Z}^{2}$ and so one expects a large set of periodic integral points within the plane region where $s_{d}$ is a good approximation of $s_{\infty}$. Interestingly, the dynamics near the boundary of this region allows for integer cycles of $h_{d}$ with long period, despite the bound of $20$ on the period of an integer cycle for the limiting map $h_{\infty}$.
@@
The organization of the paper is as follows. In Section 2 we recall the basic features of the construction of Doyle-Hyde, and in Section 3 we prove various auxiliary estimates which will allow us to understand $h_{d}$, as well as noting as an application an optimality result for Doyle-Hyde. Section 4 provides the proof of Theorem A, and Theorem B is proved in Section 5, which also includes an analysis of the limiting map $h_{\infty}.$
@@ 2. Background
In this section, we review the notion of dynamical compression and the constructions given in [doyle2022polynomialsrationalpreperiodicpoints].
@@ 2.1. Compressing polynomials
Let $K$ be a number field, $f\in K[x]$ be a polynomial and $f^{\circ n}$ be its $n$-fold composition. Recall that $\alpha\in\overline{K}$ is preperiodic if the forward orbit of $\alpha$ under $f$ is finite. Let $\operatorname{Preper}(f,K)$ denote the set of $K$-rational preperiodic points.
@@ 2.2. Discrete approximations of sine and cosine
Define the following functions of one real variable:
@@
The degree $d$ polynomial $s_{d}\in\mathbb{Q}[x]$ is integer-valued on $\mathbb{Z}+\frac{d+1}{2}$ and takes values in $\{-1,0.1\}$ for integers in the interval $[-\frac{d+1}{2},\frac{d+1}{2}].$ In [doyle2022polynomialsrationalpreperiodicpoints], Doyle and Hyde show that by applying the following transformation:
@@
The polynomial family $s_{d}$ is characterized by initial values together with the inductive relation
@@
For $d\equiv 1\mod 4,$ $s_{d}$ can be interpreted as a discrete analogue of the sine function on some finite interval; indeed, the construction of $s_{d}$ boils down to being a polynomial interpolation on the set $\left\{-\frac{d+1}{2},-\frac{d-1}{2},\cdots,\frac{d+1}{2}\right\}$ of a 6-periodic sequence $-1,-1,0,1,1,0,\cdots$. In this analogy, the $\delta$-operator is a discrete differentiation and we see that each $s_{d}$ corresponds to ±$\pm\sin$ or ±$\pm\cos$. We formalize this observation, which follows from computations in [doyle2022polynomialsrationalpreperiodicpoints]:
@@
First consider the following table:
$x$ $-3$ $-\frac{5}{2}$ $-2$ $-\frac{3}{2}$ $-1$ $-\frac{1}{2}$ 0 $\frac{1}{2}$ 1 $\frac{3}{2}$ 2 $\frac{5}{2}$ 3
$\sigma(x-3/2)$ $-1$ 0 $1$ $1$ 0 $-1$
$\sigma(x)$ 0 $-1$ $-1$ 0 1 1 0
Note that applying the central difference operator to the second row gives the first row, i.e. $\delta\sigma(x)=\sigma(x-3/2)$. Following the analogy above, $\sigma(x)$ and $\sigma(x-3/2)$ are the discrete analogues of $\sin$ and $\cos$, respectively.
From Lemma 4.1 and 4.2 of [doyle2022polynomialsrationalpreperiodicpoints] we have that $s_{d}(m-\frac{d+1}{2})=\rho(m,d)$ for integers $10\leq m\leq d+$ 1, where $\rho:\mathbb{Z}^{2}\to\mathbb{Z}$ is a doubly periodic sequence. It suffices to show that $\rho(m,d)=\sigma(m+d-2)$. We prove this using induction on $d\geq 1$. The claim is true for $d=1$, and suppose that it is true for some $d\geq 1$. Using the property $\rho(m+1,d+1)-\rho(m,d+1)=\rho(m,d)$ in Lemma 4.1 of [doyle2022polynomialsrationalpreperiodicpoints], observe that
$\rho(m+1,d+1)-\rho(m,d+1)=\sigma(m+d-2)$
$=\delta\sigma(m+d-1/2)=\sigma(m+d)-\sigma(m+d-1),$
and hence $\rho(m,d+1)-\sigma(m+d-1)$ is constant in $m$. But since both $\rho$ and $\sigma$ takes values in $\left\{-1,0,1\right\}$, we must have $\rho(m,d+1)=\sigma(m+d-1)$ as functions of $m$. Hence the claim is true for $d+1$. ∎
@@ 3. Estimates for $s_{d}$
In this section we prove a series of estimates of $s_{d}$ which allow us to study the Hénon maps of Section 4. As a consequence, we will also see in Subsection 3.3 that the dynamically compressing maps $r_{d}$ have no rational preperiodic points outside the set of integers they compress.
@@ 3.1. Local uniform convergence of $s_{d}$
In this subsection we establish the the locally uniform convergence of subsequences of $s_{d}$ to the trigonometric functions they approximate. We require careful estimates so work with $s_{d}$ as functions of a real (rather than complex) variable.
Lemma 3.1.
$(-1)^{d}s_{2d+1}(x)$ pointwise converges to the sine function $\frac{2}{\sqrt{3}}\sin\left(\frac{\pi x}{3}\right)$ as $d\to\infty$, ie.
$\sum_{j=0}^{\infty}(-1)^{j}c_{2j+1}(x)=\frac{2}{\sqrt{3}}\sin\left(\frac{\pi x}{3}\right)$
for all $x\in\mathbb{R}$.
@@
$(-1)^{j}c_{2j+1}(x)$ $=(-1)^{j}x\frac{(x+1)(x+2)\cdots(x+j)(x-1)(x-2)\cdots(x-j)}{(2j+1)!}$
$=x\frac{(1-x)(2-x)\cdots(j-x)(1+x)(2+x)\cdots(j+x)}{(\frac{3}{2})(\frac{5}{2})\cdots(\frac{2j+1}{2})}\frac{1}{4^{j}j!}$
$=x\frac{(1-x)^{(j)}(1+x)^{(j)}}{(3/2)^{(j)}}\frac{1}{4^{j}j!},$
@@
$\sum_{j=0}^{\infty}(-1)^{j}c_{2j+1}(x)=x\sum_{j=0}^{\infty}\frac{(1-x)^{(j)}(1+x)^{(j)}}{(3/2)^{(j)}}\frac{1}{4^{j}j!}=x{}_{2}F_{1}\left(1-x,1+x;\frac{3}{2};\frac{1}{4}\right),$
@@
for $|z|<1$. It is well-known that the RHS series converges absolutely on the open unit disk if $\Re(c-a-b)>-1$. We use formula 15.1.16 of [Abramowitz]:
${}_{2}F_{1}\left(a,2-a;\frac{3}{2};\sin^{2}z\right)=\frac{\sin((2a-2)z)}{(a-1)\sin(2z)}.$
Putting $a=1-x$ and $z=\pi/6$, we get the desired result of
$\sum_{j=0}^{\infty}(-1)^{j}c_{2j+1}(x)=x{}_{2}F_{1}\left(1-x,1+x;\frac{3}{2};\frac{1}{4}\right)=\frac{2}{\sqrt{3}}\sin\left(\frac{\pi x}{3}\right).$
We require a series of auxiliary bounds on $c_{d}$ and its derivative, which we prove now. For convenience, we put $c_{d}(x)$ in the same setting for both even and odd $d$ by use of the generalized binomial coefficient, which we will denote
@@
It suffices to prove that $|\tilde{c}_{d}(x)|<1/4d$ on $10\leq x\leq d-$ 1. Note $\tilde{c}_{d}(x)=0$ if $x$ is an integer between 00 and $d$. We also note that $|\tilde{c}_{d}|$ is symmetric wrt. $x=\frac{d-1}{2}$, so it suffices look at non-integers $\frac{d-1}{2}<x<d-1$. For $x<d-2$ observe that
$\left|\frac{\tilde{c}_{d}(x+1)}{\tilde{c}_{d}(x)}\right|=\frac{x+1}{d-1-x}>1.$
Hence the supremum of $|\tilde{c}_{d}|$ must be attained at $d-2+\alpha$ for some $10<\alpha<$ 1. Plugging this in and applying the AM-GM inequality, we get:
$|\tilde{c}_{d}(x)|\leq|\tilde{c}_{d}(d-2+\alpha)|$ $=\frac{1}{d!}\alpha(1-\alpha)(\alpha+1)(\alpha+2)\cdots(\alpha+d-2)$
$\leq\frac{1}{d!}\frac{1}{4}(d-1)!=\frac{1}{4d}.$
for all $0\leq x\leq d$. ∎
@@
It suffices to prove that $|\tilde{c}_{d}(x)|\leq 1/(2d(d-1))$ on $21\leq x\leq d-$ 2. By the exact same argument as the proof of Lemma 3.2, the supremum of $|\tilde{c}_{d}(x)|$ on $21\leq x\leq d-$ 2 occurs at $d-3+\alpha$ for some $10<\alpha<$ 1. Plugging this in,
$|\tilde{c}_{d}(d-3+\alpha)|=\frac{1}{d!}(1-\alpha)(2-\alpha)\alpha(\alpha+1)\cdots(\alpha+d-3).$
We have $22-\alpha<$ 2, $\alpha(1-\alpha)\leq 1/4$ and $(\alpha+1)\cdots(\alpha+d-3)<(d-2)!$. These give us the bound:
$|\tilde{c}_{d}(x)|\leq|\tilde{c}_{d}(d-3+\alpha)|<\frac{2}{4}\frac{(d-2)!}{d!}=\frac{1}{2d(d-1)}.$
Lemma 3.4.
$|c_{d}^{\prime}(x)|<\frac{\log d+3}{2d}$ for $|x|\leq\frac{d-1}{2}$.
Proof.
@@
It suffices to prove that $|\tilde{c}_{d}^{\prime}(x)|<\frac{\log d+3}{2d}$ on $10\leq x\leq d-$ 1. By the product rule, we have the following expression:
$\tilde{c}_{d}^{\prime}(x)=\frac{1}{d!}\sum_{0\leq j<d}\ \prod_{0\leq i<d,i\neq j}(x+i).$
Note that $|\tilde{c}_{d}^{\prime}(x)|\leq 1/d<(\log d+3)/2d$ for integer $x$.
Now take non-integer $N<x<N+1$ for $N=0,\cdots,d-2$. For $j<N$, note
$|\prod_{i\neq j}(x+i)|<|\prod_{i=0}^{d-1}(x+i)|\cdot\frac{1}{N-j}.$
Similarly for $j>N+1$, we have
$|\prod_{i\neq j}(x+i)|<|\prod_{i=0}^{d-1}(x+i)|\cdot\frac{1}{j-N-1}.$
For the products corresponding to $j=N$ and $j=N+1$, we obtain:
$|\prod_{i\neq N}(x+i)|+|\prod_{i\neq N+1}(x+i)|$
$<(N+1-x)\cdot(d-N-1)!(N+1)!+(x-N)\cdot(d-N-1)!(N+1)!$
$=(d-N-1)!(N+1)!.$
Applying the triangle inequality to $\tilde{c}_{d}^{\prime}(x)$ and combining all of the bounds above,
$|\tilde{c}_{d}^{\prime}(x)|<$ $\left(\sum_{j=0}^{N-1}\frac{1}{N-j}+\sum_{j=N+2}^{d-1}\frac{1}{j-N-1}\right)\left|\frac{1}{d!}\prod_{i=0}^{d-1}(x+i)\right|+\frac{(d-N-1)!(N+1)!}{d!}.$
Here observe that $\sum_{j=0}^{N-1}\frac{1}{N-j}+\sum_{j=N+2}^{d-1}\frac{1}{j-N-1}=H_{N}+H_{d-N-2}$, where $H_{n}$ denotes the $n$th harmonic number. Since $H_{n}<\log n+1$, we have
$H_{N}+H_{d-N-2}<\log(N(d-N-2))+2<2\log d+2.$
We also note $\left|\frac{1}{d!}\prod_{i=0}^{d-1}(x+i)\right|=|\tilde{c}_{d}(x)|\leq\frac{1}{4d}$ by Lemma 3.2. Finally,
$\frac{(d-N-1)!(N+1)!}{d!}=\frac{1}{\binom{d}{N+1}}\leq\frac{1}{d}.$
where its maximum is attained at $N=0$ or $N=d-2$. Combining these three bounds, we obtain the following result:
$|\tilde{c}_{d}^{\prime}(x)|<\frac{2\log d+2}{4d}+\frac{1}{d}=\frac{\log d+3}{2d}.$
Lemma 3.5.
$|c_{d}^{\prime}(x)|<\frac{\log d+3}{d(d-1)}$ for $|x|\leq\frac{d-3}{2}$.
Proof.
It suffices to prove that $|\tilde{c}_{d}^{\prime}(x)|<\frac{\log d+3}{d(d-1)}$ on $21\leq x\leq d-$ 2. We replicate the proof of Lemma 3.4, bearing in mind that there is a tighter constraint on $x$.
If $x$ is an integer between 1 and $d-2$, then
$|\tilde{c}_{d}^{\prime}(x)|=\frac{x!(d-x-1)!}{d!}=\frac{1}{d\binom{d-1}{x}}\leq\frac{1}{d(d-1)}\left(<\frac{\log d+3}{d(d-1)}\right),$
where the maximum occurs at $x=1$. Now consider non-integer $N<x<N+1$ for $N=1,\cdots,d-3$. We obtain the same bound:
$|\tilde{c}_{d}^{\prime}(x)|$ $<\left(\sum_{j=0}^{N-1}\frac{1}{N-j}+\sum_{j=N+2}^{d-1}\frac{1}{j-N-1}\right)\left|\frac{1}{d!}\prod_{i=0}^{d-1}(x+i)\right|+\frac{(d-N-1)!(N+1)!}{d!}.$
Recall that $\sum_{j=0}^{N-1}\frac{1}{N-j}+\sum_{j=N+2}^{d-1}\frac{1}{j-N-1}<2\log d+2$. We also have
$\left|\frac{1}{d!}\prod_{i=0}^{d-1}(x+i)\right|=|\tilde{c}_{d}(x)|\leq\frac{1}{2d(d-1)}$
for $|x|\leq d-1$ by Lemma 3.3. Finally, we have
$\frac{(d-N-1)!(N+1)!}{d!}=\frac{1}{\binom{d}{N+1}}\leq\frac{2}{d(d-1)}.$
where the maximum is attained at $N=1$ or $N=d-3$. As before, these bounds combine to yield
$|\tilde{c}_{d}^{\prime}(x)|<\frac{2\log d+2}{2d(d-1)}+\frac{2}{d(d-1)}=\frac{\log d+3}{d(d-1)}.$
From the four auxiliary bounds on $c_{d}$ and its derivative, we immediately obtain a stronger convergence result.
Proposition 3.6.
$(-1)^{d}s_{2d+1}(x)$ and its derivative converge locally uniformly to $\frac{2}{\sqrt{3}}\sin\left(\frac{\pi x}{3}\right)$ and its derivative, respectively, on $\mathbb{R}$.
Proof.
Recall that $(-1)^{d}s_{2d+1}(x)=\sum_{j=0}^{d}(-1)^{j}c_{2j+1}(x)$. It suffices to show uniform convergence of $(-1)^{d}s_{2d+1}(x)$ and its derivative on $[-N,N]$ for all integers $N\geq 1$. On that interval, we have asymptotic bounds:
$|c_{2j+1}(x)|=O(j^{-2}),\ |c^{\prime}_{2j+1}(x)|=O(j^{-2}\log j)$
for all $j\geq N+1$, by Lemmas 3.3 and 3.5. Both $\sum j^{-2}$, $\sum j^{-2}\log j$ are convergent series, thus we have uniform convergence of $(-1)^{d}s_{2d+1}(x)$ and $(-1)^{d}s_{2d+1}^{\prime}(x)$ on $X$ by the Weierstrass-M test.
Given a sequence $\left\{f_{n}\right\}$ of continuously differentiable real-valued functions on compact $[a,b]\subset\mathbb{R}$, recall the following fact: if $f_{n}^{\prime}$ and $f_{n}$ converge uniformly on $[a,b]$, then $f=\lim f_{n}$ is differentiable and $f^{\prime}$ is the limit of $f_{n}^{\prime}$. Using this, we conclude that $(-1)^{d}s_{2d+1}^{\prime}(x)$ converges uniformly to the derivative of $\frac{2}{\sqrt{3}}\sin(\pi x/3)$ on $X$. ∎
A similar analysis yields the analogous result in even degree, which we state without proof details.
Proposition 3.7.
$(-1)^{d}s_{2d}(x)$ and its derivative converge locally uniformly to $\frac{2}{\sqrt{3}}\cos\left(\frac{\pi x}{3}\right)$ and its derivative, respectively, on $\mathbb{R}$.
3.2. Tail behaviour of $s_{d}$
We use the results of the previous subsection to prove a lower bound on $s_{d}$ on the integers $\left\{\frac{d+7}{2},\frac{d+9}{2},\cdots\right\}$. This result will be useful in proving the boundedness of $K(r_{d})$ to understand the preperiodic points of $r_{d}$, and will also be used in Section 4 when analyzing the periodic points of the Hénon map $h_{d}$.
Lemma 3.8.
Proof.
We proceed by induction on $d$. The claim is true for $s_{3}(x)=\frac{x^{3}-7x}{6}$. Now take $d\geq 4$ and suppose claim is true for some $d-1$. Take $x=\frac{d+7+2k}{2}$ with integer $k\geq 0$, then by $s_{d-1}=\delta s_{d}$ we have
$s_{d}(x)$ $=s_{d}\left(\frac{d+5}{2}\right)+s_{d-1}\left(\frac{d+6}{2}\right)+s_{d-1}\left(\frac{d+8}{2}\right)+\cdots+s_{d-1}\left(\frac{d+6+2k}{2}\right)$
$\geq s_{d}\left(\frac{d+5}{2}\right)+\frac{3}{2}\left((d+6)+(d+8)+\cdots+(d+6+2k)\right)$
$=s_{d}\left(\frac{d+5}{2}\right)+\frac{3}{2}(k+1)(d+6+k)$
from the induction hypothesis. Since $s_{d}(\frac{d+5}{2})\geq d+1$ by Lemma 4.2(5) of [doyle2022polynomialsrationalpreperiodicpoints], we have
$s_{d}(x)-3x$ $\geq d+1+\frac{3}{2}(k+1)(d+6+k)-\frac{3}{2}(d+7+2k)$
$=d-\frac{1}{2}+\frac{3}{2}k(k+d+5)$
$\geq\frac{9}{2}+\frac{3}{2}k(k+9)>0$
for all $k\geq 0$ and $d\geq 4$. Therefore the claim is true for $d$. ∎
Now we will show that $s_{d}$ is strictly increasing on $[\frac{d+3}{2},\infty)$; this is achieved by first showing that $s_{d}$ is strictly increasing on $[\frac{d+3}{2},\frac{d+5}{2}]$ and then applying a double induction argument.
@@ 3.2. Tail behaviour of $s_{d}$
For $d\leq 9$, this is easily verified by numerical calculation, so it suffices to prove that $s_{d}^{\prime}(x)>0$ on $[\frac{d+3}{2},\frac{d+5}{2}]$ for all $d\geq 10$.
@@
$<\frac{\log d+3}{2d}+\sum_{k=4}^{\infty}\frac{\log(d+2k)+3}{(d+2k)(d+2k-1)}=:E_{d}$
by Lemmas 3.4 and 3.5. Since $\frac{\log x}{x}$ is decreasing for $x\geq e$ and numerical calculation gives $E_{d}<1$ for $d=10$, it follows that $E_{d}<1$ for all $d\geq 10$. Hence it suffices to show that
$c_{d+2}^{\prime}(x)-c_{d+4}^{\prime}(x)>\frac{2\pi}{3\sqrt{3}}+1$
@@
$\tilde{c}_{d+2}^{\prime}(x)-\tilde{c}_{d+4}^{\prime}(x+1)>\frac{2\pi}{3\sqrt{3}}+1$

### v2

@@ Abstract.
In this paper we study the dynamical uniform boundedness conjecture in the context of Hénon maps. Building on work of Doyle and Hyde on polynomial maps in one variable, we construct for each odd integer $d\geq 3$ a Hénon map of degree $d$ defined over $\mathbb{Q}$ with at least $(d-4)^{2}$ integral periodic points and an integral cycle of length at least $\frac{8d+10}{3}$. This provides a quadratic lower bound on any conjectural uniform bound for periodic rational points of Hénon maps, and answers in the affirmative a question of Ingram concerning the existence of many rational periodic points and long rational cycles for Hénon maps.
@@ 1. Introduction
In this article, we study the polynomials constructed by Doyle and Hyde and use them to construct an explicit family of Hénon maps $\{h_{d}(x,y)\}$ over $\mathbb{Q}$ with degree $d\rightarrow\infty$ and a large number of rational periodic points and long rational cycles. This answers in the affirmative the first two parts of the following question of Ingram [henon_openqs]:
Question (Ingram).
Over a number field $K$, is it possible to construct infinite families of generalized Hénon maps of algebraic degree $d$, and $K$-rational cycles of length at least $d+3$? Can one construct maps with $N_{d}$ periodic points, where $N_{d}-d\rightarrow\infty$, or even $N_{d}/d\rightarrow\infty$ as $d\rightarrow\infty$?
@@
A direct computation shows (see section 5.2) that any point of $\mathbb{C}^{2}$ with integer coordinates is periodic under iteration of the limiting Hénon map $h_{\infty}(x,y)=(y,-x+s_{\infty}(y)$, with period $n\in\{1,4,5,6,12,20\}$ determined by the class of $(x,y)$ modulo $6$.
Though the proof of Theorem A is not asymptotic, each map $h_{d}$ preserves the lattice $\mathbb{Z}^{2}$ and so one expects a large set of periodic integral points within the plane region where $s_{d}$ is a good approximation of $s_{\infty}$. Interestingly, the dynamics near the boundary of this region allows for integer cycles of $h_{d}$ with long period, despite the bound of $20$ on the period of an integer cycle for the limiting map $h_{\infty}$: see Figure 1.
Figure 1. Periodic integer cycles for $h_{d}$ with $d=43$ plotted in $\mathbb{Z}^{2}$ and colored by period length. There are a large number of periodic points with cycle length at most $20$, but the boundary of this region admits a cycle of length $118$.
@@
The organization of the paper is as follows. In Section 2 we introduce the polynomials $s_{d}$ used to construct the Hénon maps $h_{d}$, and in Section 3 we prove various auxiliary estimates which will allow us to understand $h_{d}$, as well as noting as an application an optimality result for Doyle-Hyde. Section 4 provides the proof of Theorem A, and Theorem B is proved in Section 5, which also includes an analysis of the limiting map $h_{\infty}.$
@@ 2. Background
In this section, we introduce the polynomials $s_{d}$ as integer-valued approximations of discrete trigonometric functions. We also review the notion of dynamical compression and the constructions given in [doyle2022polynomialsrationalpreperiodicpoints].
@@ 2.1. Compressing polynomials
Let $K$ be a number field, $f\in K[x]$ be a polynomial and $f^{n}:=f^{\circ n}$ be its $n$-fold composition. Recall that $\alpha\in\overline{K}$ is preperiodic if the forward orbit of $\alpha$ under $f$ is finite. Let $\operatorname{Preper}(f,K)$ denote the set of preperiodic points in $K$.
@@ 2.2. Discrete approximations of sine and cosine
Define the following functions of one variable:
@@
The degree $d$ polynomial $s_{d}\in\mathbb{Q}[x]$ is integer-valued on $\mathbb{Z}+\frac{d+1}{2}$ and takes values in $\{-1,0,1\}$ for integers in the interval $[-\frac{d+1}{2},\frac{d+1}{2}].$ In [doyle2022polynomialsrationalpreperiodicpoints], Doyle and Hyde show that by applying the following transformation:
@@
The polynomial family $s_{d}$ is alternatively characterized by initial values together with the inductive relation
@@
For $d\equiv 1\mod 4,$ $s_{d}$ can be interpreted as a discrete analogue of the sine function on some finite interval; indeed, the construction of $s_{d}$ boils down to being a polynomial interpolation on the set $\left\{-\frac{d+1}{2},-\frac{d-1}{2},\cdots,\frac{d+1}{2}\right\}$ of a 6-periodic sequence $-1,-1,0,1,1,0,\cdots$. Applying the discrete $\delta$, see that each $s_{d}$ corresponds to ±$\pm\sin$ or ±$\pm\cos$. We formalize this observation, which summarizes the computation of subsections 4.2 and 4.4 in [doyle2022polynomialsrationalpreperiodicpoints]:
@@
This proof rephrases the definition of $s_{d}$, which was given in terms of the doubly periodic sequence $\rho(m,d)$ in [doyle2022polynomialsrationalpreperiodicpoints], in a more intuitive way. Consider the following table, where each row is 6-periodic and each column is 4-periodic:
$d$ $m$cdots $-3$ $-\frac{5}{2}$ $-2$ $-\frac{3}{2}$ $-1$ $-\frac{1}{2}$ 0 $\frac{1}{2}$ 1 $\frac{3}{2}$ 2 $\frac{5}{2}$ 3 $\cdots$
1 0 $-1$ $-1$ 0 1 1 0
2 $1$ 0 $-1$ $-1$ 0 $1$
3 0 $1$ $1$ 0 $-1$ $-1$ 0
4 $-1$ 0 $1$ $1$ 0 $-1$
5 0 $1$ $1$ 0 $-1$ $-1$ 0
Note that applying the central difference operator $\delta$ to the $(d+1)$-th row gives the $d$-th row. Unpacking the definition given in [doyle2022polynomialsrationalpreperiodicpoints], it is straightforward to check that $\rho(m,d)=\sigma(m+d-2)$. Since $s_{d}$ is defined by the interpolation $s_{d}(m-\frac{d+1}{2})=\rho(m,d)$ on integers $10\leq m\leq d+$ 1, and the $d$-th row is given by values of the function $\sigma(m+\frac{3}{2}(d-1))$, it follows that $s_{d}$ interpolates the centremost $d+2$ values of the $d$-th row above.
@@ 3. Estimates for $s_{d}$
In this section we prove a series of estimates of $s_{d}$ which allow us to study the Hénon maps of Section 4. As an application of these estimates to one-dimensional dynamics, we will also see in Subsection 3.3 that the dynamically compressing maps $r_{d}$ have no rational preperiodic points outside the set of integers they compress; that is, the bounds given for the explicit examples of Section 4 of [doyle2022polynomialsrationalpreperiodicpoints] are sharp.
@@ 3.1. Local uniform convergence of $s_{d}$
In this subsection we establish the local uniform convergence of subsequences of $s_{d}$ to the trigonometric functions they approximate.
Proposition 3.1.
$(-1)^{d}s_{2d+1}(z)$ locally uniformly converges to the sine function $\frac{2}{\sqrt{3}}\sin\left(\frac{\pi z}{3}\right)$ as $d\to\infty$ on the complex plane; that is,
$\sum_{j=0}^{\infty}(-1)^{j}c_{2j+1}(z)=\frac{2}{\sqrt{3}}\sin\left(\frac{\pi z}{3}\right)$
locally uniformly on $z\in\mathbb{C}$.
@@
$(-1)^{j}c_{2j+1}(z)$ $=(-1)^{j}z\frac{(z+1)(z+2)\cdots(z+j)(z-1)(z-2)\cdots(z-j)}{(2j+1)!}$
$=z\frac{(1-z)(2-z)\cdots(j-z)(1+z)(2+z)\cdots(j+z)}{(\frac{3}{2})(\frac{5}{2})\cdots(\frac{2j+1}{2})}\frac{1}{4^{j}j!}$
$=z\frac{(1-z)^{(j)}(1+z)^{(j)}}{(3/2)^{(j)}}\frac{1}{4^{j}j!},$
@@
$\sum_{j=0}^{\infty}(-1)^{j}c_{2j+1}(z)=z\sum_{j=0}^{\infty}\frac{(1-z)^{(j)}(1+z)^{(j)}}{(3/2)^{(j)}}\frac{1}{4^{j}j!}=z{}_{2}F_{1}\left(1-z,1+z;\frac{3}{2};\frac{1}{4}\right),$
@@
for $c\not\in\mathbb{Z}_{\leq 0}$ and $|z|<1$. Note that the ratios between consecutive terms converge to
$\frac{(a+n)(b+n)z}{(c+n)n}\to z,$
which implies that the series converges uniformly on compact subsets of
$\left\{(a,b,c,z)\in\mathbb{C}^{4}:c\not\in\mathbb{Z}_{\leq 0},\ |z|<1\right\}.$
We use formula 15.1.16 of [Abramowitz]:
${}_{2}F_{1}\left(a,2-a;\frac{3}{2};\sin^{2}w\right)=\frac{\sin((2a-2)w)}{(a-1)\sin(2w)}.$
Putting $a=1-z$ and $w=\pi/6$, we get the desired result of
$\sum_{j=0}^{\infty}(-1)^{j}c_{2j+1}(z)=z{}_{2}F_{1}\left(1-z,1+z;\frac{3}{2};\frac{1}{4}\right)=\frac{2}{\sqrt{3}}\sin\left(\frac{\pi z}{3}\right),$
where the convergence is locally uniform in $z\in\mathbb{C}$. ∎
Since all $s_{d}$ and the sine function are holomorphic, it follows from standard results in complex analysis that the derivative of $(-1)^{d}s_{2d+1}(z)$ also locally converges to the derivative of the sine wave. A similar argument shows that $(-1)^{d}s_{2d}(z)$ and its derivative locally uniformly converge to $\frac{2}{\sqrt{3}}\cos(\frac{\pi z}{3})$ and its derivative on the complex plane.
3.2. Tail behaviour of $s_{d}$
We now work with $s_{d}$ as functions of a real (rather than complex) variable and show that $s_{d}$ is strictly increasing on the real interval $[\frac{d+3}{2},\infty)$; this is achieved by first showing that $s_{d}$ is strictly increasing on $[\frac{d+3}{2},\frac{d+5}{2}]$ and then applying a double induction argument.
For our analysis, we need some explicit auxiliary bounds on the derivative of $c_{d}$ on the real line. For convenience, we put $c_{d}(x)$ in the same setting for both even and odd $d$ by use of the generalized binomial coefficient, which we denote
@@ 3.2. Tail behaviour of $s_{d}$
We prove that $|\tilde{c}_{d}^{\prime}(x)|\leq 1$ on $10\leq x\leq d-$ 1. By the product rule, we have the following expression:
$\tilde{c}_{d}^{\prime}(x)=\frac{1}{d!}\sum_{0\leq j<d}\ \prod_{0\leq i<d,i\neq j}(x-i).$
By the triangle inequality,
$|\tilde{c}_{d}^{\prime}(x)|\leq\frac{1}{d!}\sum_{0\leq j<d}\ \prod_{0\leq i<d,i\neq j}|x-i|\leq\frac{1}{d!}\sum_{0\leq j<d}(d-1)!=1.$
@@
We prove that $|\tilde{c}_{d}^{\prime}(x)|\leq\frac{6+4\log d}{d(d-1)}$ on $21\leq x\leq d-$ 2. By continuity of $\tilde{c}_{d}^{\prime}$, it suffices to consider non-integer $x$. We carry on from the expression above:
$|\tilde{c}_{d}^{\prime}(x)|\leq\frac{1}{d!}\sum_{0\leq j<d}\ \prod_{0\leq i<d,i\neq j}|x-i|.$
Fix an integer $21\leq N<d-$ 2 and take $N<x<N+1$. Note that for $j<N$,
$\prod_{i\neq j}|x-i|\leq\frac{(N+1)!(d-N-1)!}{x-j}\leq\frac{(N+1)!(d-N-1)!}{N-j}.$
Similarly for $j>N+1$,
$\prod_{i\neq j}|x-i|\leq\frac{(N+1)!(d-N-1)!}{j-x}\leq\frac{(N+1)!(d-N-1)!}{j-N-1}.$
For $j=N$ and $N+1$, we simply get
$\sum_{j=N,N+1}\ \prod_{i\neq j}|x-i|\leq((x-N)+(N+1-x))\cdot\prod_{i\neq N,N+1}|x-i|\leq(N+1)!(d-N-1)!.$
Combining these bounds, we obtain:
$|\tilde{c}_{d}^{\prime}(x)|$ $\leq\frac{1}{d!}(N+1)!(d-N-1)!(1+H_{N-1}+H_{d-N})$
$\leq\frac{1+2H_{d}}{{}_{d}C_{N+1}}\leq\frac{1+2(\log d+1)}{d(d-1)/2}=\frac{6+4\log d}{d(d-1)},$
where $H_{n}=1+1/2+\cdots+1/n<\log n+1$. ∎
@@
For $d\leq 44$, this is verified by numerical calculation, so it suffices to prove that $s_{d}^{\prime}(x)>0$ on $[\frac{d+3}{2},\frac{d+5}{2}]$ for all $d>44$.
@@
$<1+\sum_{k=4}^{\infty}\frac{6+4\log(d+2k)}{(d+2k)(d+2k-1)}=:E_{d}$
by Lemmas 3.2 and 3.3. Since each of the summands are decreasing in $d$ and numerical calculation gives $E_{d}<2$ for $d=44$, it follows that $E_{d}<2$ for all $d>44$. Hence it suffices to show that
$c_{d+2}^{\prime}(x)-c_{d+4}^{\prime}(x)>\frac{2\pi}{3\sqrt{3}}+2$
@@
$\tilde{c}_{d+2}^{\prime}(x)-\tilde{c}_{d+4}^{\prime}(x+1)>\frac{2\pi}{3\sqrt{3}}+2$